ULTRAFAST EXCITED STATE DYNAMICS OF NON-LINEAR OPTICAL CHROMOPHORES PROBED BY FS/PS-CARS

\underline{BENJAMIN D. PRINCE},MIKHAIL N. SLIPCHENKO, BETH M. PRINCE, ALEX J. BLOM, and HANS U. STAUFFER, Department of Chemistry, Iowa State University, Ames, Iowa, 50011-3111.

The development of a time-resolved coherent anti-Stokes Raman scattering (CARS) variant has been applied to the excited state vibrational dynamics of model non-linear optical chromophores in the condensed phase. This variant, termed fs/ps-CARS combines two initial femtosecond laser pulses (ω_{1} and ω_{2}) generating vibrational population and a third picosecond narrowband probe to integrate the molecular response. The resulting signal contains reasonable spectral width (ω_{3} bandwidth limited with a typical linewidth of ω_{3} in the 6-15 cm^{−1}regime) while maintaining the femtosecond time resolution of the first two pulses. This technique proves amenable to a detailed study of the vibrational dynamics of excited state molecules by the addition of a femtosecond excitation pulse (ω_{ex}) with a total time resolution of about 150 fs. As a demonstration, the excited state dynamics, known to internally convert on sub-picosecond timescales, of two model chromophores (para-nitroaniline and N,N-dimethyl-para-nitroaniline) are followed.